[Brahms-l] Shift report 20040228 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Feb 28 2004 - 23:58:54 EST
------------------------------------------------------------
Date: Sat Feb 28 23:58:53 2004, Shift: 16:00-24:00
Supervisor: YKL, with: 
Log book#15 pages 68 - 68

Shift Summary:
------------------------------------------------------------
1. TQ tripped and the repair took 3 hours.
Otherwise we had good beams.

2. MRS @40 degrees B2255
   FS @8 degrees B1723

run 10660  Tr2=145K Tr3=183K
    10661       13       33
    10664      101      137
    10665       76      142
    10666       43      108
    10667 continuing

Run 10662 is TPC pedestals 
    10663 is non-TPC pedestals

3. HV channels C1B01 through C1B16 for C1
are individually tweeked to default
voltages and they are stable for several
hours
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Feb 28 23:59:16 2004

This archive was generated by hypermail 2.1.8 : Sat Feb 28 2004 - 23:59:37 EST